Problem 3 (Quiz). Recall that an American Roulette wheel has 38 possible outcomes, 18 black, 18 red and 2 green. The numbers go from 1 to 36, and the green outcomes are numbered 0 and 00. Someone places İO bets on 1st 12, that is, that person wins if the outcome is any number between 1 and 12, included. 1. What is the probability of losing 8 out of the 10 total bets?
Answer #1

Binomial distribution: P(X) = nCx px qn-x

p = P(losing one bet) = 1 - P(winning one bet)

= 1 - 12/38

= 13/19

q = 1 - p = 6/19

n = 8

P(losing 8 out of 10 bets) = P(X = 8)

= 10C8 x (13/19)8 x (6/19)2

= 0.2155